mathslover (mathslover):

AB is any diameter of the circle \(x^2 + y^2 + 2x + 2y = 0 \) and CD is any diameter of circle \(x^2 + y^2 + x + y - 4 = 0 \) such that A, B , C and D are con cyclic. The locus of center of the circle passing through A,B,C,D is ...

mathslover (mathslover):

@ganeshie8

mathslover (mathslover):

It's not a big deal. Here is the solution, but, I want someone to explain me that.

mathslover (mathslover):

I'm not able to understand that from where did the radical axis come? I know that they have used that as a trick, but if there is a possible proof, then I would love to see it.

ganeshie8 (ganeshie8):

when circles are tangent to each other, radical axis is the common tangent i still don't see how this will be the center of circle passing through A,B,C,D

mathslover (mathslover):

Exactly! :'(

mathslover (mathslover):

Now, let us mark the diameters and all using geogebra. May be that will help us a bit.

mathslover (mathslover):

\(x^2 + y^2 + 2x + 2y = 0 \) Center = (-1,-1) Radius = \(\sqrt{2}\) x^2 + y^2 + x + y - 4 =0 Center = \((-1/2 , -1/2 )\) Radius = \(3/\sqrt{2}\)

mathslover (mathslover):

Looks like this.

mathslover (mathslover):

Well, I will see what I can do with diameters now.

myininaya (myininaya):

what does concylic mean?

myininaya (myininaya):

they lie on a common circle ?

mathslover (mathslover):

Yeah!

ganeshie8 (ganeshie8):

I am not able to visualize how that straight line gives the locus of center of circles passing through points A,B,C,D

ganeshie8 (ganeshie8):

*radical axis

ganeshie8 (ganeshie8):

can you draw atleast one circle that contains the diameter endpoints A,B,C,D ?

mathslover (mathslover):

Yeah, one minute. I will try.

mathslover (mathslover):

Interesting.

ganeshie8 (ganeshie8):

also draw the locus line from solution

ganeshie8 (ganeshie8):

does the center of that circle fall on that locus line ?

mathslover (mathslover):

I wonder how to draw the center of that circle.

mathslover (mathslover):

I will try, one sec.

mathslover (mathslover):

http://web.geogebra.org/app/?id=qEuAL9Nx Well, can you do that on geogebra?

mathslover (mathslover):

I've shared the link here. I am not sure how will we be able to mark the center of that circle.

ganeshie8 (ganeshie8):

it is easy to eyeball actually it is 1 km away from the locus line in solution

mathslover (mathslover):

Aw... sorry, I've never used geogebra that much. No idea what is eyeball :/

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

the centers of the circles that pass through points A,B,C,D must lie on that radical axis, yes ?

ganeshie8 (ganeshie8):

as per your solution ^

mathslover (mathslover):

Yes!

mathslover (mathslover):

Oops... light cut-off! Its raining here... :'( I need to sign out now bhaiya! Sorry. Can we discuss 'bout this tomorrow?

mathslover (mathslover):

(if you don't mind).

ganeshie8 (ganeshie8):

yea sure :)

OpenStudy (anonymous):

lala